### v2

# Upsample or Upweight? Balanced Training on Heavily Imbalanced Datasets

Abstract
Data availability across domains often follows a long-tail distribution: a few domains have abundant data, while most face data scarcity. This imbalance poses challenges in training language models uniformly across all domains. In our study, we focus on multilingual settings, where data sizes vary significantly between high- and low-resource languages. Common strategies to address this include upsampling low-resource languages (Temperature Sampling) or upweighting their loss (Scalarization). Although often considered equivalent, this assumption has not been proven, which motivates our study. Through both theoretical and empirical analysis, we identify the conditions under which these approaches are equivalent and when they diverge. Specifically, we demonstrate that these two methods are equivalent under full gradient descent, but this equivalence breaks down with stochastic gradient descent. Empirically, we observe that Temperature Sampling converges more quickly but is prone to overfitting. We argue that this faster convergence is likely due to the lower variance in gradient estimations, as shown theoretically. Based on these insights, we propose Cooldown, a strategy that reduces sampling temperature during training, accelerating convergence without overfitting to low-resource languages. Our method is competitive with existing data re-weighting and offers computational efficiency.
Upsample or Upweight?
Balanced Training on Heavily Imbalanced Datasets
Tianjian Li, Haoran Xu, Weiting Tan Kenton Murray, Daniel Khashabi Center for Language and Speech Processing Johns Hopkins University tli104@jhu.edu
1 Introduction
The internet is full of information, ranging from common knowledge like famous landmarks and popular recipes to rare details such as local folklore and specialized scientific theories. Some data are widely available, while others are much harder to collect (Kandpal et al., 2023). Data availability across different domains is often long-tailed (Feldman, 2020; Feldman and Zhang, 2020), with few domains having an abundance of data and many rare domains where data is scarce. However, the standard language model training objective weighs each training instance equally, paying little attention to domains that suffer from data scarcity. The heavy mismatch in dataset sizes creates substantial challenges in training language models to be competent in all domains.
Cooldown reduces overfitting compared to upsampling (temp=5)
Cooldown converges $\sim$5$\times$ faster compared to proportional sampling (temp=1).
Figure 1: Validation loss by training iteration of a low-resource language pair (En-Ro) in multilingual machine translation. Proportional sampling leads to underfitting the low-resource direction. Using a high temperature (oversampling LRLs) leads to overfitting the low-resource direction. Employing a high temperature at the beginning and then decreasing the temperature (Cooldown) gets the advantage of fast convergence without overfitting.
Specifically, we focus on language modeling on a natural divide of domains with a heavy mismatch: different languages in multilingual language modeling. As a result of this mismatch, language models are often trained on corpora with an overwhelming amount of English and other high-resource languages (HRLs) and very few amounts of data for low-resource languages (LRLs) (Koehn and Knowles, 2017; Conneau et al., 2020; Xue et al., 2021). For example, the multilingual C4 corpus (Xue et al., 2021) contains 2733 billion English tokens but only 1 billion Swahili tokens. Uniformly sampling from the large, combined dataset would result in the language model optimized heavily towards performance on HRLs (e.g. English), sacrificing performance on LRLs (e.g. Swahili).
Two methods are often employed to address domain mismatches: Scalarization and Temperature Sampling. Scalarization adjusts the losses for individual domains by re-weighting them under uniform sampling (Zhou et al., 2021; Choi et al., 2023). In this case, we assign a larger weight to LRLs to emphasize their importance. Temperature Sampling, weights each training instance uniformly and handles the mismatch by over-sampling HRLs and/or down-sampling LRLs. (Aharoni et al., 2019; Wang et al., 2020b; Chung et al., 2023; Xue et al., 2021). Intuitively, Scalarization modifies the loss while Temperature Sampling modifies the dataset. Scalarization and Temperature Sampling are widely regarded as equivalent. Choi et al. (2023) denotes “we follow convention and implement Scalarization via proportional sampling”. Xie et al. (2023) and Fan et al. (2024) implement sampling probabilities by multiplying losses with per-domain re-normalized weights. The underlying assumption is that Temperature Sampling and Scalarization are equivalent, and we can use them interchangeably. However, to the best of our knowledge, this assumed equivalence has neither been proven nor disproven.
We closely investigate this assumed equivalency theoretically (§3.1). Specifically, we prove that although they are equivalent in full gradient descent (Theorem 1), Temperature Sampling induces lower variance in the context of stochastic gradient descent (Theorem 2). Moreover, the variance induced by scalarization increases as the approximated temperature increases or the domain distribution’s skewness increases (Theorem 3). Based on our theoretical results and connecting to the literature on lower variance between stochastic gradients accelerates convergence Sutskever et al. (2013), we make the following hypothesis:
Hypothesis 1.
Temperature Sampling (TS) converges much faster than Scalarization (S) at higher temperatures or on heavily imbalanced domain distributions.
We empirically verify our hypothesis (§3.2) and find that Temperature Sampling does converge faster but is more prone to overfitting. We identify in TS, the temperature controls the speed of convergence and hence can be used as a control knob to adjust the convergence speed. We thus propose Cooldown: to use a large temperature initially for fast convergence, then decrease the temperature to prevent overfitting to the LRLs. Figure 1 illustrates the effectiveness of Cooldown, which significantly accelerates convergence on the LRL due to a high temperature (aggressive upsampling of LRLs) at the beginning of training and reduces overfitting to the LRL due to the lowering the temperature during training.
To sum up, our contribution is two-fold:
  • We inspect Scalarization and Temperature Sampling both theoretically and empirically (§3). Contrary to existing work that uses them interchangeably, we found that Temperature Sampling converges faster due to a lower variance in stochastic gradient estimation.
  • Motivated by our findings, we propose Cooldown, a method to adjust the sampling temperature during training on unbalanced datasets. We show the effectiveness of Cooldown on two tasks: multilingual machine translation and multilingual language modeling (§4).
2 Preliminaries
2.1 Notations and Task Description
We consider a model is trained on a collection of data $\mathcal{D}=\{x\}_{i=1}^{N}$ from $K$ domains $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup...\cup\mathcal{D}_{K}$. Here, “domain" refers to sources (Books, Wikipedia, code) for general language modeling or different languages (English, French, Swahili) in multilingual language modeling. The total training loss $\mathcal{J}(D)$ is the sum of the losses of each example $\mathcal{L}(x)$.
$\mathcal{J}(D)=\sum_{x\in\mathcal{D}}\mathcal{L}(x).$
Scalarization (S)
Naive aggregation often results in imbalanced performance across domains when high-resource domains dominate the aggregated loss. Scalarization solves this issue by assigning weights $\{w_{i}\}_{i=1}^{K}$ to each domain and aggregates the weighted sum of individual losses:
$\mathcal{L}_{\textrm{S}}=\sum_{x\in\mathcal{D}}w_{f(x)}\mathcal{L}(x),$
where $f:\mathcal{D}\rightarrow[K]$ maps a training example to the index of its domain. Scalarization balances the loss by assigning a higher weight to harder or low-resource domains.
Temperature Sampling (TS)
Instead of assigning weights to losses, we can also sample more frequently from the low-resource domain to achieve balanced training. Temperature sampling achieves this by adjusting the probabilities of selecting instances from different domains based on their sizes. The sampling probability vector of each domain $p$ is given by:
$\forall i\in\{1,2,...,K\}:\;p(\mathcal{D}_{i};\tau)=\frac{|\mathcal{D}_{i}|^{\frac{1}{\tau}}}{\sum_{j=1}^{K}|\mathcal{D}_{j}|^{\frac{1}{\tau}}},$
where $\tau$ is the sampling temperature, a hyperparameter controlling the sampling weights. $\tau=1$ means that we are sampling proportional to the sizes of each domain. As we increase $\tau$, we increase the sampling probability of low-resource domains. The loss for Temperature Sampling is:
$\mathcal{L}_{\textrm{TS}}=\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p}\left[\sum_{x\in\mathcal{D}_{i}}\mathcal{L}(x)\right].$
The common understanding is that Temperature Sampling is equivalent to Scalarization (Choi et al., 2023; Kurin et al., 2022; Xin et al., 2022), and we can use them interchangeably (Xie et al., 2023). In the next subsection, we will formalize this statement and show that these two are mathematically equivalent in full gradient descent. We will then show that they are not equivalent under stochastic gradient descent.
3 Temperature Sampling V.S. Scalarization
3.1 Theoretical Analysis
We formalize the equivalence of Scalarization and weighted sampling under full-gradient descent (Theorem 1) and show that Scalarization induces a larger variance between the mini-batch losses (Theorem 2). Furthermore, when using Scalarization to approximate Temperature Sampling, the variance increases as the temperature rises (Theorem 3).
Theorem 1 (Equivalency under Gradient Descent).
For any sampling temperature $\tau$, there exists a set of weights $\textbf{w}=\{w_{1},w_{2},...,w_{K}\}$ for the Scalarization loss such that this loss is equivalent to the Temperature Sampling loss, both computed based on the whole data $\mathcal{D}$.
Proof.
For all $i\in\{1,2,3,...,K\}$, let:
$w_{i}=\frac{\sum_{j=1}^{K}|D_{j}|}{|D_{i}|}\cdot\frac{|\mathcal{D}_{i}|^{\frac{1}{\tau}}}{\sum_{j=1}^{K}|\mathcal{D}_{j}|^{\frac{1}{\tau}}}=\frac{p(\mathcal{D}_{i};\tau)}{p(\mathcal{D}_{i};1)},$
then
$\mathcal{L}_{\textrm{S}}$ $=\mathop{\mathbb{E}_{x\sim D}}\left[w_{f(x)}\mathcal{L}(x)\right]$
$=\sum_{x\in\mathcal{D}}\frac{p(\mathcal{D}_{f(x)};1)}{|\mathcal{D}_{f(x)}|}w_{f(x)}\mathcal{L}(x)$
$=\sum_{i=1}^{K}p(\mathcal{D}_{i};\tau)\sum_{x\in\mathcal{D}_{i}}\frac{1}{|\mathcal{D}_{i}|}\mathcal{L}(x)$
$=\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}\left[\sum_{x\in\mathcal{D}_{i}}\frac{1}{|\mathcal{D}_{i}|}\mathcal{L}(x)\right]$
$=\mathbb{E}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}[\mathbb{E}_{x\sim\mathcal{D}_{i}}[\mathcal{L}(x)]]=\mathcal{L}_{\textrm{TS}}.$
Intuitively, this suggests that in the context of full gradient descent, the loss remains the same whether you multiply the loss of a single data point by 2 (S) or duplicate the data point (TS). We will then show that Scalarization induces a larger variance in stochastic gradient estimation (Robbins, 1951) compared to Temperature Sampling.
Corollary 1.1.
Denote the total gradient as $\nabla\mathcal{L}_{S}=\nabla\mathcal{L}_{TS}$, and $\nabla\mathcal{L}(x)$ is the gradient with respect to a single datapoint $x$. We denote the stochastic gradient under Scalarization as $\nabla\mathcal{L}_{\textrm{S}}(x)=\{\nabla w_{f(x)}\mathcal{L}(x)|x\sim\mathcal{D}\}$, which means the stochastic gradient $\nabla w_{f(x)}\mathcal{L}(x)$ when $x$ is uniformly drawn from the dataset $\mathcal{D}$. Similarly, we denote the stochastic gradient under Temperature Sampling as $\nabla\mathcal{L}_{\textrm{TS}}(x)=\{\nabla\mathcal{L}(x)|x\sim\mathcal{D}_{i},\mathcal{D}_{i}\sim p(\mathcal{D}_{i};\tau)\}$. Then both $\nabla\mathcal{L}_{\textrm{S}}(x)$ and $\nabla\mathcal{L}_{\textrm{TS}}(x)$ are unbiased estimates of the total gradient.
Proof.
By Theorem 1, $\nabla\mathcal{L}_{s}=\mathop{\mathbb{E}}_{x\sim\mathcal{D}}\left[w_{f(x)}\nabla\mathcal{L}(x)\right]$, and $\nabla\mathcal{L}_{TS}=\mathbb{E}_{\mathcal{D}_{i}\sim p(\cdot;\tau)}[\mathbb{E}_{x\sim\mathcal{D}_{i}}[\nabla\mathcal{L}(x)]]$.
We conclude that
$\mathop{\mathbb{E}_{\mathcal{D}_{i}\sim p(\mathcal{D}_{i};\tau)}\left[\mathbb{E}_{x\sim\mathcal{D}_{i}}\left[\nabla\mathcal{L}(x)\right]\right]}$
$=\mathbb{E}_{x\sim\mathcal{D}}$ $[w_{f(x)}\nabla\mathcal{L}(x)].$
Theorem 2 (Scalarization induces larger variance under Stochastic Gradient Descent).
Using the same notation in Corollary 1.1, we have $\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))\geq\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$.
We provide a detailed proof in Appendix B.
Implication of Theorem 2
Scalarization induces larger variance connects with the literature on variance reduction in stochastic gradient estimation (Sutskever et al., 2013; Kingma and Ba, 2015) accelerates the convergence of SGD (Robbins, 1951). We thus hypothesize that Temperature Sampling will converge faster than Scalarization with the set of weights that makes them mathematically equivalent under full gradient descent.
Theorem 3 (Scalarization induces larger variance when approximating higher temperatures).
The difference $\Delta=\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))-\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$ monotonically increases when $\tau\geq 1$.
Figure 2 illustrates the variance by increasing temperature and skewness of the distribution. We defer details on how we constructed Figure 2 and the full proof of Theorem 3 to Appendix C.
Figure 2: Variance of Scalarization $\sum_{i}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}$ by sampling temperature $\tau$. A large temperature or a skewed distribution of $\mathcal{D}$ induces a much larger variance for Scalarization. Distributions $\mathcal{D}_{i}\propto\frac{1}{i^{\alpha}}$. See Appendix C for details of the experiment setup.
Implication of Theorem 3
The fact that the induced variance of Scalarization increases as the approximated temperature increases implies that Temperature Sampling converges much faster than Scalarization at higher temperatures, which we empirically verify in the next section.
3.2 Empirical Evidence
[…]
Language Pairs Training Validation
WMT 15 En-Fr 40M 4503
WMT 22 En-Zh 55M 3418
WMT 22 En-Cs 56M 2082
WMT 16 En-Ro 600K 1999
[…]
Figure 3 shows the variance between mini-batch gradients for Temperature Sampling and Scalarization. Figure 4 shows the validation loss curves by training iterations.
Figure 3: Gradient Norm distributions on En-{Cs, Ro} for sampling and Scalarization. Empirical results show that Scalarization induces a larger variance (2.25 >>> 0.62) between mini-batch gradients compared to Temperature Sampling.
Figure 4: Validation loss by training iteration for En-{Cs, Ro} (first row) and En-{Fr, Ro} (second row). Temperature Sampling (dashed) converges faster compared to Scalarization (solid), leading to better performance on both the HRL and the LRL.
We found that temperature sampling converges faster than scalarization at all temperatures we experimented with. This supports our theory that Temperature Sampling lowers the variance in gradient estimation (Theorem 2), thus speeding up convergence. The larger the temperature, the larger the gap between the convergence speed of Temperature Sampling v.s. Scalarization. We conclude that when there is a heavy mismatch in sizes, it is better to use Temperature Sampling with a large temperature, confirming the intuition that when the LRL has very little data, we want to upsample it more aggressively. We summarize our findings below:
Large Temperature Sampling is prone to overfitting
In our En - {Fr, Ro} experiments, we observed that the model overfits the low-resource direction (Ro) when using a large temperature ($\tau=3,5$). However, the high-resource direction has not yet converged when this overfitting occurs; therefore, continuing to train in both directions would lead to severe overfitting. This indicates that we need to pair strong regularization on the LRL (e.g. early stopping) with a large temperature, which motivates our temperature scheduling method Cooldown, which uses a large temperature during the beginning to speed up training and then decreases the temperature to prevent overfitting on the LRL.
Temperature Sampling is equivalent to Scalarization given enough compute
We found that Temperature Sampling (dashed in Figure 4) always converges faster compared to weighting the losses of individual directions (solid in Figure 4), but they eventually converge to the same validation loss given enough training iterations, which corresponds to their equivalency under full gradient descent (Theorem 1). This means both Scalarization and Temperature Sampling can effectively balance multiple languages when given a large enough compute.
4 Cooldown: Balanced Training for Heavily Imbalanced Datasets
Based on the theoretical analysis in §3.1 and the empirical results in §3.2, we conclude that Temperature Sampling with a large temperature converges faster than Scalarization with equivalent weights but is more prone to overfitting on the LRL when using a large temperature. We thus hypothesize that we can employ a large temperature during the beginning of training to speed up convergence and then decrease the temperature to prevent overfitting on low-resource directions.
Our proposed method:
We design a simple temperature scheduling method: Cooldown which starts with a high temperature (aggressive upsampling of LRLs) and then lowers the temperature to $\tau=1$ (proportional sampling) at a fixed iteration to prevent overfitting. We describe our experiment setup at §4.1, report results at §4.2, and discuss our findings at §4.3.
4.1 Setup
Models, Datasets, and Hyper-parameters
We experiment on two setups: multilingual machine translation and multilingual language modeling, both suffering severe mismatch in dataset sizes. For our machine translation experiments, we use the standard encoder-decoder Transformer (Vaswani et al., 2017) architecture implemented in fairseq (Ott et al., 2019). We select 8 distinct languages from the opus-100 dataset (Zhang et al., 2020) and train a one-to-many translation where the source language is English. Detailed Languages and their respective sizes can be found in Table 4. For our multilingual language modeling experiments, we use a decoder-only Transformer model from Huggingface (Wolf et al., 2020) and select 4 linguistically diverse languages with varying amounts of data from the mC4 (Xue et al., 2021) dataset. The statistics are in Table 2.
For our machine translation experiments, we use $\tau=5$ for the first 30k of training iterations and use $\tau=1$ for the second 30k. For our language modeling experiments, we use $\tau=5$ for the first 50k training iterations and $\tau=1$ for the second 50k. More detailed hyper-parameters are in Appendix D.
Baselines
[…]
4.2 Main Results
Languages Tokens
mT5 weights
$\tau$ = 3.33
EN - English 2733B 5.67%
IT - Italian 162B 2.43%
ZH - Chinese 39B 1.67%
SW - Swahili 1B 0.5%
Table 2: Dataset statistics of our selected subset of C4. mT5 (Xue et al., 2021) effectively uses an sampling temperature of $\tau=3.33$ to oversample LRLs.
EN IT ZH SW
Cooldown 2.59 2.63 2.56 2.94
Dense Update 1 3.17 2.89 2.85 3.11
Dense Update 2 3.31 3.02 3.35 3.16
Table 3: Validation loss on MC4 subset (lower is better) for different update schedules in Figure 5. Sparse updates perform better than dense updates.
Machine Translation
Table 4 shows applying Cooldown on training a multilingual machine translation model. Compared to proportional sampling ($\tau=1$), Cooldown is able to greatly improve the mid and low-resource languages (+0.7 and 3.1 BLEU) while minimally sacrificing the performance of high-resource languages (-0.2 BLEU). Compared to static up-sampling $\tau=5$, Cooldown matches the performance of mid- and low-resource languages while improving the performance of high-resource languages by 1.5 BLEU. Our method also outperforms the performance of Unimax and Order Matters scheduling while being easier to implement. Furthermore, Cooldown is able to match the performance of DoReMi without having to train multiple proxy models.
[…]
# of Parallel Sentences 1M 1M 294K 519K 102K 538K 1M 83K 76K $>$1M $<$500K
Static Temperature Sampling
$\tau=1$ (Proportional) 38.9 13.1 58 28.9 41.9 17.1 32.8 22.4 10.8 28.3 30.4
$\tau=5$ 36.9 12.2 60.9 28.3 46.3 16.8 30.7 26.8 9.4 26.6 32.4
[…]
Dynamic Temperature Sampling
Unimax (Chung et al., 2023) 37.1 12.2 60.7 28.4 45.9 16.5 30.8 26.2 9.4 26.7 32.1
Order Matters (Choi et al., 2023) 37.1 12.2 60.9 28.2 46.1 16.7 30.8 26.7 9.4 26.7 32.3
Cooldown (Ours) 38.7 13.2 60.1 28.7 47.4 17.3 32.2 26 11 28.1 32.4
With Proxy Model Training
DoReMi (Xie et al., 2023) 37.3 13.1 60.4 28.4 46.3 17.1 29.8 23.0 10.8 26.7 32.4
Table 4: SacreBLEU scores (higher is better) on a chosen subset of OPUS-100 with a mixture of high (1M), mid (500K - 1M), and low ($<$500K) resource languages. The best performance is bolded. Scores that are close (within 1 BLEU) of the best performance are colored in green. Scores lower than the best for more than 1.5 BLEU are highlighted in red. Our simple scheduling method outperforms various static and dynamic Temperature Sampling methods. We improve the performance on LRLs without sacrificing much performance on HRLs.
Multilingual Language Modeling
We also experimented with the general language modeling task on multiple languages on selected languages on the multilingual C4 (mC4) dataset (Xue et al., 2021). We report the validation loss in table 5. Our results echo the findings in our machine translation experiments: Cooldown matches the performance on the only HRL English (EN) but outperforms other baselines on all three other languages.
Method EN 2733B IT 162B ZH 39B SW 1B
Static Temperature Sampling
$\tau=1$ (Proportional) 2.67 3.10 4.23 4.09
$\tau=5$ 2.91 3.01 3.14 3.12
[…]
Dynamic Temperature Sampling
Order Matters (Choi et al., 2023) 2.77 2.76 3.23 3.26
Unimax (Chung et al., 2023) 2.98 2.91 2.85 3.06
Cooldown (Ours) 2.75 2.63 2.56 2.94
With Proxy Model Training
DoReMi (Xie et al., 2023) 2.89 2.63 2.51 2.89
Table 5: Dev loss on selected mC4 subset (lower is better). Cooldown achieves the best performance on all three LRLs (IT, ZH, and SW). On the only HRL English, Cooldown is only behind proportional sampling ($\tau$=1), which is heavily optimized towards performance on English by sacrificing performance on all other languages.
4.3 Findings / Ablations
In this section, we revisit two design choices in dynamic Temperature Sampling: 1) Increasing v.s. Decreasing the temperature during training, and 2) Dense v.s. Sparse updates of the temperatures. We highlight our results below:
Figure 5: Sampling temperature schedules.
Increasing the temperature is better than decreasing the temperature.
Unimax (Chung et al., 2023) upsamples the LRLs during the beginning of training while Order Matters (Choi et al., 2023) upsamples LRLs at the end of training. To resolve this conflict, we compare an increasing temperature schedule (1 for the first 15k training iterations and 5 for the second 15k training iterations; “1-5") with a decreasing temperature schedule (5 for the first 15k iterations and 1 for the second half; “5-1"), using the same En-{Zh, Ro} machine translation data in §3.2. Figure 7 illustrates that a decreasing schedule (dashed) converges faster and results in better performance on the LRL compared to an increasing schedule (solid), with minimal sacrifice on the HRL. This means that upsampling the LRLs during the beginning of training performs better.
Curse of Granularity
We concluded that a decreasing schedule generally leads to faster convergence and better overall performance. Furthermore, we compare various fine-grained decreasing schedules that perform dense decreasing of the sampling temperature with our sparse update (5 for the first 50k, 1 for the second 50k training iterations) using the subset in mC4 described in Table 2. Figure 5 illustrates the schedules we compare, and 3 shows the validation loss of each decreasing schedule. Fine-grained online reweighting method yields worse performance, echoing the findings in Fan et al. (2024) and Xie et al. (2023). [^3] Both Xie et al. (2023) and Fan et al. (2024) approximate sampling probabilities using Scalarization when performing dense updates, which could also be the reason why dense update underperforms in their experiments.
Figure 6: Comparison between an increasing (solid) Temperature Sampling schedule and a decreasing (dashed) schedule in multilingual machine translation En-{Zh, Ro}. A decreasing temperature schedule outperforms an increasing one on the LRL with minimal sacrifice on the HRL.
5 Related Works
Gradient-based methods for Multi-Task Optimization
Training a multilingual model can be seen as Multi-Task Learning (MTL; Caruana (1997)), where each language is a single task. Gradient-based methods aim to reduce the discrepancy in directions of conflicting gradients in different tasks: PCGrad (Yu et al., 2020) aims to project the gradient of a task onto the orthogonal plane of the gradients of the other task, which explicitly aligns the gradient of different tasks. Gradient Vaccine (Wang et al., 2021) points out that using PCGrad to optimize multilingual pre-training relies on the false assumption that all languages are equally related and proposes to align the similarity of the gradient to the relatedness of the two languages. Another line of work (Wang et al., 2020a, b; Kreutzer et al., 2021) uses gradient similarity as the reward to train policy that decides the sampling probabilities for each domain in a Reinforcement Learning setting. Fan et al. (2024) utilizes gradient similarity between domains to design a temperature schedule for balancing multiple domains in training language models. However, recent studies point out that such gradient-based techniques do not yield significant improvement compared to a weighted sum of individual task losses (Scalarization) (Kurin et al., 2022; Xin et al., 2022; Royer et al., 2023). Our results also echo the findings of Zhai et al. (2023), where they find loss reweighting (Scalarization) underperforms standard training. Our work provides a possible explanation that Scalarization induces a larger variance in gradients.
Loss-based methods for Multi-Task Learning
Another line of work utilizes the loss instead of the gradients per task for optimizing MTL models. An intuitive method is to put more weight on the task with the highest loss. In statistical learning, Distributionally Robust Optimization (DRO) methods (Ben-Tal et al., 2011; Duchi and Namkoong, 2021; Hashimoto et al., 2018; Sagawa* et al., 2020) instead of minimizing the sum of all losses, minimize the loss of the worst-performing subgroup to balance performance across all subgroups. Oren et al. (2019) and Xie et al. (2023) apply DRO to multi-domain language modeling to minimize the loss of a set of worse-performing domains. Similarly, Zhou et al. (2021) applies DRO to multilingual machine translation by minimizing the loss of a set of worse-performing translation pairs. One drawback of DRO-based methods is that they require training a proxy model on the dataset, which hinders their applicability on large-scale datasets. Cooldown can be seen as an efficient approximation of DRO methods by upsampling the worse-performing LRL and shifting the focus to the HRL once the LRL is sufficiently trained. Unlike DRO-based methods, Cooldown do does not require training proxy models (Liu et al., 2021) and dense updates on domain weights/sampling probabilities.
6 Conclusion
We examined two common balancing methods for multi-domain language modeling with data imbalances: Scalarization and Temperature Sampling. Although both yield the same loss, Temperature Sampling converges faster but risks overfitting. To mitigate this, we propose Cooldown, a variant that adjusts temperatures during training to maintain fast convergence while reducing overfitting.
Limitations
We discuss the limitations of our study here.
  1. 1.
    Impact of data mixture on downstream performance Studies have pointed out that the data mixtures in different domains impact downstream performance (Gururangan et al., 2020; Albalak et al., 2023; Fan et al., 2024). Our work only focuses on the impact of different temperature schedules on pre-training validation performance. Although the effect of optimizing pre-training mixtures across different languages on downstream performances has not been fully concluded, works have shown that a lower pre-train validation loss generally leads to better downstream performance (Xie et al., 2023; Du et al., 2024).
  2. 2.
    Difference between multi-lingual and multi-domain language modeling Existing work on mono-lingual language modeling (Chowdhery et al., 2023; Du et al., 2022; Xie et al., 2023; Oren et al., 2019; Fan et al., 2024; Longpre et al., 2023) maps all data from the same source (e.g. Wikipedia, Web, Books) to a single domain. Such a mapping ignores the subdomains within each source. In our work, we focused on the multilingual setup because (a) There exists a severe data size mismatch and (b) there is a clear and natural definition of “domain" — the different languages, and we expect the sampling to have a larger impact because of this mismatch. Even though we only conducted experiments on a multilingual setup, our theoretical analysis applies to all setups with heavy dataset size mismatches.
  3. 3.
    Finding the optimal temperature schedule It requires a large amount of compute to thoroughly study the scaling laws of pre-training language models under different temperatures (Ye et al., 2024) and to search for the optimal static sampling temperature $\tau$ for a given dataset (Chen et al., 2023), let along dynamic temperature scheduling. The optimal temperature depends not just on the size of the dataset but also on the “difficulty" of the dataset. Therefore, existing research (Chen et al., 2023; Xie et al., 2023; Oren et al., 2019; Zhou et al., 2021; Liu et al., 2024; Dubey et al., 2024) relies on training a proxy model on the dataset probe domain difficulty to determine optimal weights. Our work, instead, proposes a heuristic that decreases the temperature during training that does not rely on any proxy training, but we did not exhaustively test all the decreasing schedules. We leave finding optimal temperature schedules without preliminary training for future work.
Ethical Considerations
One application of our study is to balance high- and low-resource languages. We aim to mitigate biases that favor high-resource languages. However, this approach also raises risks, such as amplifying existing biases in limited and potentially skewed low-resource corpora. Furthermore, improved models could be misused to spread misinformation or infringe on privacy, especially in communities less equipped to counter such impacts. Thus, while Cooldown  promotes linguistic diversity, it requires careful monitoring to ensure it is used ethically.
Acknowledgements
This work is supported by a generous gift the Allen Institute for AI and partly by ONR grant (N00014-24-1-2089). We are grateful to Jingyu Zhang for his insightful feedback throughout this project. GPU machines for conducting experiments were provided by the DSAI cluster.
Supplemental Material
Appendix A Additional Related Works
Multilingual Interference
Finding out the reasons and solutions for negative interference in Multilingual Neural Machine Translation (Johnson et al., 2017; Aharoni et al., 2019) has been an active research area for the past decade. Yet, while previous studies (Wang et al., 2021) find that negative interference mainly occurs between different language families, recent studies (Shaham et al., 2023) have demonstrated that negative inference does not happen between languages of different families. The interference emerges because of the mismatch in the amount of data for different translation directions. Real-world translation data suffers from a heavy mismatch of data size in different directions, ranging from less than 100K to over 100M (NLLB Team et al., 2022). In our work, we show that this heavy mismatch in data size results in low-resource languages being under-trained.
To mitigate interference caused by dataset sizes, Aharoni et al. (2019) and Xue et al. (2021) propose to up-sample low-resource languages, which often results in the model overfitting on the LRLs while underfitting HRLs. Huang et al. (2022) proposes to distill the model from earlier checkpoints with the LRLs that have not overfit with the current model to regularize the training of LRLs. Huang et al. (2023) proposes to distill between a model trained with a low sampling temperature and a model trained with a high sampling temperature. Unimax (Chung et al., 2023) proposes to first uniformly sample from all languages until an LRL dataset has been seen by the model for a fixed amount of repetitions; then, we remove the LRL from training. Order-Matters (Choi et al., 2023) proposes the opposite of Unimax (Chung et al., 2023), to first only train on the HRL and add in the LRL after a fixed iteration. Our work shows that the Unimax style of decreasing the temperature works better and proposes a simple alternative that does not require tracking how many times a model has seen an individual LRL dataset.
Appendix B Proof of Theorem 2
Theorem 2 (Scalarization induces larger variance under Stochastic Gradient Descent) Using the same notation in Corollary 1.1, we have $\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))\geq\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$.
We first proof a required lemma:
Lemma 2.1 Let $\mathcal{D}=\{\mathcal{D}_{1},...,\mathcal{D}_{K}\}$, $\forall j\in\{1,...,K\}$, let $p(\mathcal{D}_{j},\tau)=\frac{\mathcal{D}_{j}^{\frac{1}{\tau}}}{\sum_{k}\mathcal{D}_{k}^{\frac{1}{\tau}}}$, then $\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}\geq 1$ holds true when $\forall j,\mathcal{D}_{j}\geq 0$.
Proof.
We substitute $x_{j}=\mathcal{D}_{j}^{\frac{1}{\tau}}$, then:
$\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)$
By the Cauchy-Schwartz inequality, we have:
$\left(\sum_{i}x_{i}^{\tau/2}\cdot x_{i}^{({2-\tau})/2}\right)^{2}\leq\left(\sum_{i}x_{i}^{\tau}\right)\left(\sum_{i}x_{i}^{2-\tau}\right),$
which simplifies to:
$\left(\sum_{i}x_{i}\right)^{2}\leq\left(\sum_{i}x_{i}^{\tau}\right)\left(\sum_{i}x_{i}^{2-\tau}\right),$
which implies that:
$1\leq\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)=\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}.$
Armed with Lemma 2.1, we come back to the proof of Theorem 2.
Proof.
Since $\nabla\mathcal{L}_{\textrm{S}}(x)$ and $\nabla\mathcal{L}_{\textrm{TS}}(x)$ are unbiased estimates of the total gradient, we only need to show that the expectation of the squared gradient is larger for the stochastic gradient under Scalarization.
$\underbrace{\mathop{\mathbb{E}}_{x\sim\mathcal{D}}[(w_{f(x)}\nabla\mathcal{L}(x))^{2}]}_{\text{Scalarization}}$ $=\sum_{i=1}^{K}p(\mathcal{D}_{i};1)w^{2}_{f(x)}\nabla\mathcal{L}^{2}(x)$
$\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}\nabla\mathcal{L}^{2}(x)$ $\geq\nabla\mathcal{L}^{2}(x)\sum_{i=1}^{K}p(\mathcal{D}_{i};\tau)\nabla\mathcal{L}^{2}(x)$
$=\underbrace{\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p(\mathcal{D}_{i};\tau)}\left[\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[\nabla\mathcal{L}^{2}(x)]\right]}_{\text{Temperature Sampling}}.$
Appendix C Proof of Theorem 3 and Construction of Figure 2
Theorem 3 (Scalarization induces larger variance when approximating higher temperatures)
The difference in variance $\Delta=\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))-\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$ is non-decreasing when $\tau\geq 1$.
Proof.
From the proof of Theorem 2, we know that the difference in variance $\Delta$ can be quantified by $\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}$. We substitute $x_{j}=\mathcal{D}_{j}^{\frac{1}{\tau}}$, then:
$\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right).$
Let $F(\tau)=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)$ be a function of $\tau$. Taking its derivative with respect to $\tau$:
$\frac{dF(\tau)}{d\tau}=F(\tau)\times\left[\left(\sum_{i}\frac{x_{i}^{\tau}\log x_{i}}{\sum_{j}x_{j}^{\tau}}\right)-\left(\sum_{i}\frac{x_{i}^{2-\tau}\log x_{i}}{\sum_{j}x_{j}^{2-\tau}}\right)\right].$
Taking the derivative of the term $\sum_{i}\frac{x_{i}^{\tau}\log x_{i}}{\sum_{j}x_{j}^{\tau}}$ with respect to $\tau$, we get:
$\frac{d}{d\tau}\sum_{i}\left(\frac{x_{i}^{\tau}\log x_{i}}{\sum_{j}x_{j}^{\tau}}\right)=\frac{\sum_{j}x_{j}^{\tau}\sum_{i}x_{i}^{\tau}(\log x_{i})^{2}-(\sum_{i}x_{i}^{\tau}\log x_{i})^{2}}{(\sum_{i}x_{i}^{\tau})^{2}},$
which is the variance of $\log x_{i}$ under the probability distributions $\frac{x^{\tau}_{i}}{\sum_{i}x^{\tau}_{i}}$. Similarly, the derivative of $\sum_{i}\frac{x_{i}^{2-\tau}\log x_{i}}{\sum_{j}x_{j}^{2-\tau}}$ with respect to $\tau$ is the negative variance of $\log x_{i}$ under distribution $\frac{x_{i}^{2-\tau}}{\sum_{i}x_{i}^{2-\tau}}$. Since variances are always non-negative, we conclude that the following difference:
$\left[\left(\sum_{i}\frac{x_{i}^{\tau}\log x_{i}}{\sum_{j}x_{j}^{\tau}}\right)-\left(\sum_{i}\frac{x_{i}^{2-\tau}\log x_{i}}{\sum_{j}x_{j}^{2-\tau}}\right)\right],$
is always non-negative. By Lemma 2.1, we know that $F(\tau)\geq 1$. Therefore, the derivative $\frac{dF(\tau)}{d\tau}$ is always non-negative when $\tau\geq 1$, meaning that $F(\tau)$ is non-decreasing when $\tau\geq 1$.
Furthermore, when $\tau$ is strictly larger than 1 and not all $\mathcal{D}$ are equal, $F(\tau)$ monotonically increases with $\tau$. Showing that approximating a larger temperature using Scalarization induces a larger variance than approximating smaller temperatures. ∎
Construction of Figure 2
We plot the function $F(\tau)=\sum_{i}\frac{p(D_{i};\tau)^{2}}{p(D_{i};1)}$ against $\tau$ by starting with a uniform distribution and progressively increasing its skewness to resemble Zipf distributions. Specifically, we generate distributions $D_{i}\propto\frac{1}{i^{\alpha}}$ for various exponents $\alpha$ (ranging from 0 to higher values), where $i$ denotes the rank of each element. For each distribution, we compute the normalized probabilities $p(D_{i};\tau)=\frac{D_{i}^{1/\tau}}{\sum_{j}D_{j}^{1/\tau}}$ across a range of $\tau$ values. This approach allows us to analyze how increasing the skewness of the distribution influences the behavior of $F(\tau)$ as a function of $\tau$.
Appendix D Detailed Hyper-Parameters
We provide a comprehensive list of the hyper-parameters we used in this appendix section: §3.2 - Table 6.
Hyper-parameter Value Hyper-parameter Value
Arch wmt_en_de_big Label smoothing 0.1
Optimizer adam Adam epsilon 1e-06
Adam betas "(0.9, 0.98)" Learning rate scheduler inverse_sqrt
Learning rate 0.0005 Warmup updates 4000
Validate interval updates 1000 Dropout 0.1
Attention dropout 0.1 Weight decay 0.0
Max tokens 32768 Update frequency 8
Max source positions 256 Max target positions 256
Table 6: Detailed Hyper-parameters for experiments in §3.2. We use the fairseq (Ott et al., 2019) implementation.
Hyper-parameter Value Hyper-parameter Value
Arch iwslt_de_en Label smoothing 0.1
Optimizer adam Adam epsilon 1e-06
Adam betas "(0.9, 0.98)" Learning rate scheduler inverse_sqrt
Learning rate 0.0005 Warmup updates 4000
Validate interval updates 1000 Dropout 0.1
Attention dropout 0.1 Weight decay 0.0
Max tokens 16384 Update frequency 4
Max source positions 256 Max target positions 256
Table 7: Detailed Hyper-parameters for Machine Translation experiments in §4 on our selected subset of opus-100 (Zhang et al., 2020). We use the fairseq (Ott et al., 2019) implementation. Differences against Table 6 are in red.
Appendix E Additional Results on Scalarization V.S. Temperature Sampling
Figure 7: Validation loss by training iteration for En-{Zh, Ro}. Temperature Sampling (Dashed) converges much faster than Scalarization (Solid), especially at higher temperatures.
(a) validation loss: low resource languages
(b) validation loss: high resource languages
Figure 8: Validation loss by gradient updates on the low-resource and high-resource language (left) jointly trained on the same model. Adjusting the sampling temperature has little impact on the high-resource language but a high impact on the low-resource language.